INTERSTELLAR N_{2} IN THE DIFFUSE ISM TOWARD 20 AQUILAE

\underline{DAVID C. KNAUTH}, Department of Physics and Astronomy, Northwestern University, Evanston, IL 60208; B-G ANDERSSON, S. R. McCANDLISS, H. W. MOOS, Department of Physics and Astronomy, The Johns Hopkins University, Baltimore, MD 21218.

Molecular nitrogen (N_{2}) is the most abundant molecule in the Earth’s atmosphere and in the less chemically-processed atmosphere of Titan. N_{2} is also of considerable interest for studies of the interstellar medium because both models of steady-state gas-phase chemistry (Viala 1986) and millimeter wave observations of N_{2}H^{+} (Womack, Ziurys, & Wyckoff 1992) predict that N_{2} should be the dominant nitrogen-bearing molecule in interstellar space. We previously presented the first detection of interstellar N_{2} toward HD 124314 with an observed fractional abundance of N_{2}/H_{2} = x(N_{2}) = 3.3 × 10^{−7} (Knauth et al. 2004). Here we report on the second detection of interstellar N_{2} based on Far Ultraviolet Spectroscopic Explorer (F USE) observations of the moderately-reddened star 20 Aquilae (Knauth et al. 2005) with x(N_{2}) ∼ 4.7 × 10^{−8}. This fractional abundance is surprisingly low given that 20 Aql has a factor-of-4 higher CN abundance than that found toward HD 124314. Is this possible anti-correlation caused by differences in the predictions of diffuse and dark cloud chemistry? Recent work (Herbst 2004, private communication; Roueff 2005) show that a single homogeneous cloud chemical model cannot reproduce the observations. Further observations of interstellar N_{2} and other nitrogen-bearing species (e.g., NH) toward lines of sight with different physical conditions are required to understand the implications for interstellar nitrogen chemistry.